# Whose Boat Does it Float? Improving Personalization in Preference Tuning via Inferred User Personas

### v1

@@ 3.4 Personas Help Describe Preference Data
We mainly use LLM personas for personalization (section 4), but they can also help describe trends in preference datasets. User preferences are implicit, but if LLMs articulate reasons behind these preferences, we can understand differences between chosen and rejected outputs Hoyle et al. (2023). To do this, we first form sets $\mathcal{S}_{\texttt{C}}$ and $\mathcal{S}_{\texttt{R}}$ with chosen and rejected personas from our three best LLMs: L-405B, Opus, and GPT-4o. We then find salient words in each set via $\mathcal{P}(w\,|\,\mathcal{S})$: the probability a word $w\in\mathcal{S}_{\texttt{C}}\cup\mathcal{S}_{\texttt{R}}$ appears given $w$ is in the chosen or rejected set $\mathcal{S}$.
Table 1 has words with the highest $\mathcal{P}(w\,|\,\mathcal{S})$ appearing 10+ times. [^4] Low-frequency words have inflated saliency scores (e.g. appearing only once in a rejected output yields a score of one). On BeaverTails, annotators choose detailed outputs (“multiple”, “meticulous”) and reject shorter ones (“to-the-point”, “concise”), explaining verbosity bias Zheng et al. (2024). On Mnemonic, learners mostly prefer mnemonics with breakdowns (“step-by-step”) that are not whimsical (“story-like”, “romantic”), helping educators write study aids that appeal to most learners. Thus, PI is also a useful content analysis tool for preference data, informing model designers and practitioners.
@@ 4 Persona Tailoring Setup
Personas from PI are accurate (section 3), so we use them to train more personalized models that use personas and prompts as inputs to produce tailored responses. We run PI with L-405B, our best open weight LLM, to add personas to preference data, and train L-8B on this new data for Persona Tailoring (PT):
@@
Our one-time augmentation strategy is similar to knowledge distillation Gou et al. (2021): a larger teacher LLM boosts personalization in a smaller student LLM, improving efficiency for long-term deployment. This section gives our datasets (section 4.1), strategies (section 4.2) and metrics (section 4.3) used for PT.
@@ 4.1 Datasets
BeaverTails Anthropic HHH Mnemonic
In test sets, using a persona from gold outputs $r_{\texttt{C}}$ or $r_{\texttt{R}}$ may be unrealistic, as it can leak signals in $r_{\texttt{C}}$ and $r_{\texttt{R}}$. Thus, for a test set prompt $p$, we retrieve a training example $\mathcal{E}$ with the most similar prompt Wang et al. (2022) via ColBERT Santhanam et al. (2022), and use personas from $\mathcal{E}$. We refer to personas from $\mathcal{E}$ as $\mathcal{P}_{\texttt{retr}}$, and from $r_{\texttt{C}}$/$r_{\texttt{R}}$ as $\mathcal{P}_{\texttt{gold}}$; we use both for thorough testing (examples in Appendix A.1). LLM personas are imperfect proxies for real needs, so users also write personas in section 5.4.
@@ 4.2 Personalization Techniques
1) Few-shot (Brown et al., 2020, FS) has 5 examples to produce $r_{\texttt{C}}$ from $p$ and $\mathcal{P}_{\texttt{C}}$ with the template:
@@
3) Direct preference optimization (Rafailov et al., 2024, DPO) tunes a model $\pi$ from the SFT model $\pi_{0}$, using $p$ and $\mathcal{P}_{\texttt{C}}$ as inputs $x$. DPO increases the probability gap for output sequences $r_{\texttt{C}}$ and $r_{\texttt{R}}$:
@@
While $\mathcal{P}_{\texttt{R}}$ explains who may prefer $r_{\texttt{R}}$ (section 3.1), this is not causal: users in $\mathcal{P}_{\texttt{R}}$ may not be best satisfied by $r_{\texttt{R}}$ Jin et al. (2021)—the goal of PT. Empirically, PT does not benefit much when trained on rejected signals $\mathcal{P}_{\texttt{R}}$ and $r_{\texttt{R}}$, (Appendix A.8), indicating $r_{\texttt{R}}$ has lower average quality than $r_{\texttt{C}}$. Thus, we train our PT models just on $r_{\texttt{C}}$ and $\mathcal{P}_{\texttt{C}}$. However, since $\mathcal{P}_{\texttt{R}}$ is valid, it can be used in inference; in section 5.4, we show $\textsc{PT}_{\textsc{dpo}}$ supports needs in $\mathcal{P}_{\texttt{R}}$, unlike DPO.
We use greedy decoding, but on Anthropic, this sometimes yields repetitive, non-terminating texts (20% of times), as some outputs used to train models have this repetition. We filter such outputs for fair evaluation, but use the full test set in Appendix A.11—reflecting use cases when filtering is infeasible; our results remain strong on the full test set.
@@ 4.3 Metrics
To compare outputs from varied personalized models, we use a common method of model win-rate Liu et al. (2023) judged by Prometheus Kim et al. (2024). It judges the better model output on: 1) Response Quality: answering the prompt; and 2) Personalization: tailoring to the persona; these assess how well models attend to both PT inputs. For position bias, we judge outputs in both orders, only noting a winner if an output is picked twice, otherwise a tie. For the three judgment types (win/loss/tie), Prometheus has $66\%$ human agreement on (1) in three out-of-domain datasets Kim et al. (2024), the best open-weight judge, and $62\%$ agreement with two authors on (2) (Appendix A.4). Prometheus is imperfect, so users also assess responses in section 5.4.
There are quality/personalization tradeoffs: tailored outputs are specific and appeal to fewer users Chakraborty et al. (2024), harming judged quality. To capture this, we create $\Delta\textsc{PQ}$ to measure the mean gain in both metrics Demšar (2006). Formally, to see if a new model $\pi_{test}$ bests a baseline model $\pi_{base}$, we query Prometheus for win/tie/loss rates of $\pi_{test}$ vs $\pi_{base}$ on personalization ($p_{test}$, $p_{tie}$, $p_{base}$) and quality ($q_{test}$, $q_{tie}$, $q_{base}$). $\Delta\textsc{PQ}$ finds the mean improvement of $\pi_{test}$ vs $\pi_{base}$ on both metrics, ignoring ties, relative to a 50/50 random judge:
$p_{win}$ $=\frac{p_{test}}{p_{base}+p_{test}},\quad q_{win}=\frac{q_{test}}{q_{base}+q_{test}},$
$\Delta\textsc{PQ}$ $=\frac{1}{2}\left(\frac{p_{win}-0.5}{0.5}+\frac{q_{win}-0.5}{0.5}\right).$
If $\Delta\textsc{PQ}>0$, $\pi_{test}$ bests $\pi_{base}$ by giving more personalized and higher-quality outputs, or improves in one metric with minimal reductions in the other.
@@ 5 Evaluating Persona-Tailored Responses
We now compare Persona Tailoring (PT) to models trained on standard preference data [^6] Preference data is used to train small, domain-specific models Ouyang et al. (2022), so we omit large, general models (ChatGPT) that cannot be preference-tuned for domains., showing PT improves personalization based on LLM (section 5.1, section 5.2, section 5.3) and human judgments (section 5.4, section 5.5).
@@ 5.1 Persona Tailoring Aids Personalization
We first show the net benefits of PT, confirming it enhances personalization while maintaining quality versus models tuned on standard preference data that are not personalized. PT always improves in personalization via FS, SFT, and DPO with typically minor response quality losses, shown via large $\Delta\textsc{PQ}$ (Table 2). Thus, researchers seeking to improve personalization should elicit why users prefer responses in data collection. Further, if preference data has already been curated, abduction via PI is a simple but effective augmentation method that enhances personalization with low impact on quality.
@@ 5.2 $\textsc{PT}_{\textsc{dpo}}$ Supercharges Personalization
We verify $\textsc{PT}_{\textsc{dpo}}$ is the best method by comparing each PT model to FS without personas. Method response lengths have large differences, so for verbosity bias Zheng et al. (2024), we judge model outputs with the same sentence count; the trend is similar without this check (Appendix A.7). Each method shows $\Delta\textsc{PQ}$ gains on Mnemonic, but SFT shows minor losses on BeaverTails (Table 3). Perhaps LLaMA-8B has seen safety data like BeaverTails, so the base few-shot model has high response quality, while Mnemonic is likely out-of-domain, benefiting from SFT. Regardless, $\textsc{PT}_{\textsc{dpo}}$ has the best $\Delta\textsc{PQ}$. Thus, alignment methods like DPO on persona-augmented preference datasets better instill personalization than alternatives like SFT/FS.
@@
BT Chosen DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 46.7/29.3/24.0 38.5/30.5/31.1 +21.3
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 42.3/29.3/28.5 34.9/33.9/31.3 +12.5
BT Reject DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 45.1/31.7/23.2 35.1/32.5/32.5 +17.9
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 51.1/25.9/23.0 35.3/32.7/32.1 +21.3
HHH Chosen DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 37.2/22.6/40.2 32.6/21.3/46.0 -10.4
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 32.6/27.9/39.5 30.4/29.0/40.6 -11.9
HHH Reject DPO+$\mathcal{P}_{\texttt{retr}}$ PT+$\mathcal{P}_{\texttt{retr}}$ 48.0/21.8/30.1 39.3/25.8/34.9 +14.4
DPO+$\mathcal{P}_{\texttt{gold}}$ PT+$\mathcal{P}_{\texttt{gold}}$ 50.8/20.7/28.5 43.0/23.4/33.6 +20.2
@@
Average DPO $\textsc{PT}_{\textsc{dpo}}$ 43.4/27.3/29.3 43.8/28.8/30.0 +11.9
@@
5.3 DPO Sans Personas is Not All You Need
Having seen $\textsc{PT}_{\textsc{dpo}}$’s benefits (section 5.1, section 5.2), we now test if training on personas is needed for personalization: could DPO trained without input personas still tailor to inputted personas in inference? This could be doable, as LLMs may generalize to unseen instructions Deb et al. (2022), and personas are instructions. To answer this, we mainly study BeaverTails and Anthropic HHH, as their prompts are first-person, so we can add personas to prompts by writing them in first person (“the user is X”$\rightarrow$“I am X”). We also test Mnemonic, but since prompts in this dataset are a single vocabulary term, models may not be able to generalize to personas as instructions; this highlights another advantage of our method, as we can instill personalization in any type of preference data. We then compare if outputs from $\textsc{PT}_{\textsc{dpo}}$ or DPO better tailor to input personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ on response quality [^7] This response quality judge uses the persona and prompt as inputs, as DPO and $\textsc{PT}_{\textsc{dpo}}$ have access to both in inference. and personalization (section 4.3). This ensures DPO and $\textsc{PT}_{\textsc{dpo}}$, which are tuned on majority chosen outputs and thus may implicitly tailor to popular needs, still aid less popular but valid needs found in $\mathcal{P}_{\texttt{R}}$ (section 3.3).
$\textsc{PT}_{\textsc{dpo}}$ bests DPO in personalization and quality in 8/10 splits, reinforcing PT’s strength (Table 4). $\textsc{PT}_{\textsc{dpo}}$ also has higher improvements than DPO on $\mathcal{P}_{\texttt{R}}$ vs $\mathcal{P}_{\texttt{C}}$ (mean $\Delta\textsc{PQ}$ of 14.5 on $\mathcal{P}_{\texttt{R}}$ vs 6.4 on $\mathcal{P}_{\texttt{C}}$), so DPO sometimes tailors to needs in $\mathcal{P}_{\texttt{C}}$, but rarely the uncommon ones in $\mathcal{P}_{\texttt{R}}$. Thus, to curate harder evaluations for personalization, researchers can consider the often ignored reasons users may prefer rejected outputs and fully challenge models.
Figure 6: Evaluation of how well LLMs aid user-given needs. PT significantly improves personalization while maintaining answerability (Dror et al., 2018, 95% bootstrapped CIs).

### v2

@@ 3.4 Personas Help Describe Preference Data
We mainly use LLM personas for personalization (section 4), but they can also help describe trends in preference datasets. User preferences are implicit, but if LLMs articulate reasons behind these preferences, we can understand differences between chosen and rejected outputs Hoyle et al. (2023). Thus, we first create sets $\mathcal{S}_{\texttt{C}}$ and $\mathcal{S}_{\texttt{R}}$ with chosen and rejected personas from our three best LLMs: L-405B, Opus, and GPT-4o. We then find salient words in each set via $\mathcal{P}(w\,|\,\mathcal{S})$: the probability a word $w\in\mathcal{S}_{\texttt{C}}\cup\mathcal{S}_{\texttt{R}}$ appears given $w$ is in the chosen or rejected set $\mathcal{S}$.
Table 1 has words with the highest $\mathcal{P}(w\,|\,\mathcal{S})$ appearing 10+ times. [^5] Low-frequency words have inflated saliency scores (e.g., appearing only once in a rejected output yields a score of one). On BeaverTails, annotators choose detailed outputs (“multiple”, “meticulous”) and reject shorter ones (“to-the-point”, “concise”), explaining verbosity bias Zheng et al. (2024a). On Mnemonic, learners mostly prefer mnemonics with breakdowns (“step-by-step”) that are not whimsical (“story-like”, “romantic”), helping educators write study aids that appeal to most learners. Hence, PI is also a useful content analysis tool for preference data, informing model designers and practitioners.
@@ 4 Persona Tailoring Setup
Personas from PI are accurate (section 3), so we use them to train more personalized models that use personas and prompts as inputs to give tailored outputs. We run PI with L-405B, our best open weight LLM, [^6] Training on GPT/Claude outputs breaks terms-of-service. to add personas to preference data, and train L-8B on this new data for Persona Tailoring (PT):
@@
Our one-time augmentation strategy resembles knowledge distillation Gou et al. (2021): a larger teacher LLM boosts personalization in a smaller student LLM, improving efficiency for long-term deployment. This section gives our datasets (section 4.1), techniques (section 4.2), and metrics (section 4.3) used for PT.
@@ 4.1 Datasets
In test sets, using a persona from gold outputs $r_{\texttt{C}}$ or $r_{\texttt{R}}$ may be unrealistic, as it can leak signals in $r_{\texttt{C}}$ and $r_{\texttt{R}}$. Thus, for a test set prompt $p$, we retrieve a training example $\mathcal{E}$ with the most similar prompt via ColBERT Wang et al. (2022); Santhanam et al. (2022), and use personas from $\mathcal{E}$. We refer to personas from $\mathcal{E}$ as $\mathcal{P}_{\texttt{retr}}$, and from $r_{\texttt{C}}$/$r_{\texttt{R}}$ as $\mathcal{P}_{\texttt{gold}}$; we use both for thorough testing (examples in Appendix A.1). LLM personas are imperfect proxies for real needs, so users also write personas in section 5.4.
@@ 4.2 Personalization Techniques
1) Few-shot prompting (Brown et al., 2020, FS) uses 5 exemplars to produce $r_{\texttt{C}}$ with the template:
@@
3) Direct preference optimization (Rafailov et al., 2024, DPO) further tunes the SFT model $\pi_{0}$ with preference data to build a better model $\pi$. Given a prompt and persona as input $x=\langle p\cdot\mathcal{P}_{\texttt{C}}\rangle$, $\pi$ increases the likelihood of generating the chosen response $r_{\texttt{C}}$ over the rejected one $r_{\texttt{R}}$ by minimizing:
@@
While $\mathcal{P}_{\texttt{R}}$ explains who may prefer $r_{\texttt{R}}$ (section 3.1), this is not causal Jin et al. (2021): users in $\mathcal{P}_{\texttt{R}}$ may not be best satisfied by $r_{\texttt{R}}$—the goal of PT. Empirically, PT does not benefit much when trained on rejected signals $\mathcal{P}_{\texttt{R}}$ and $r_{\texttt{R}}$ (Appendix A.8), indicating $r_{\texttt{R}}$ has lower average quality than $r_{\texttt{C}}$. Thus, we train our PT models just on $r_{\texttt{C}}$ and $\mathcal{P}_{\texttt{C}}$. However, since $\mathcal{P}_{\texttt{R}}$ is valid, it can be used in inference; in section 5.4, we show $\textsc{PT}_{\textsc{dpo}}$ supports needs in $\mathcal{P}_{\texttt{R}}$, unlike DPO.
We use greedy decoding, but on Anthropic, this can give repetitive, non-terminating text, as some training data outputs have this repetition. We show results on the full test set, but even when filtering these cases, our results are strong (Appendix A.11).
BeaverTails Anthropic HHH Mnemonic
@@ 4.3 Metrics
To compare outputs of personalized models, we use a common method of model win-rate Liu et al. (2023) via Prometheus Kim et al. (2024), an LLM trained to compare pairs of examples on specified criteria. We use the LLM to compare model outputs on: (1) Response Quality: answering the prompt; and (2) Personalization: tailoring to the persona; these test how well models use both inputs of PT.
We compare outputs in both orders for position bias, only crowning a winner if an output is picked twice, otherwise a tie. For win/loss/tie judgments, Prometheus has $66\%$ human agreement on (1) in 3 out-of-domain datasets Kim et al. (2024), the best open-source judge, and $62\%$ agreement with two authors on (2) (Appendix A.4). LLM judges are imperfect, so users also assess responses in section 5.4.
There are quality/personalization tradeoffs: personalized responses are more specific and appeal to fewer users Chakraborty et al. (2024), lowering judged quality. To capture this, $\Delta\textsc{PQ}$ measures the average gain in both of these metrics. Formally, to check if a new model $\pi_{\text{test}}$ bests a baseline model $\pi_{\text{base}}$, we query Prometheus for win/tie/loss rates of $\pi_{\text{test}}$ versus $\pi_{\text{base}}$ on personalization ($p_{\text{test}}$, $p_{\text{tie}}$, $p_{\text{base}}$) and quality ($q_{\text{test}}$, $q_{\text{tie}}$, $q_{\text{base}}$). $\Delta\textsc{PQ}$ finds the mean improvement of $\pi_{\text{test}}$ vs $\pi_{\text{base}}$ on both metrics, ignoring ties, compared to a 50/50 random judge:
$p_{\text{win}}$ $=\frac{p_{\text{test}}}{p_{\text{base}}+p_{\text{test}}},\quad q_{\text{win}}=\frac{q_{\text{test}}}{q_{\text{base}}+q_{\text{test}}},$
$\Delta\textsc{PQ}$ $=\frac{1}{2}\left(\frac{p_{\text{win}}-0.5}{0.5}+\frac{q_{\text{win}}-0.5}{0.5}\right).$
If $\Delta\textsc{PQ}>0$, $\pi_{\text{test}}$ bests $\pi_{\text{base}}$ by giving more personalized and higher-quality outputs, or improves in one metric with minimal reductions in the other.
@@ 5 Evaluating Persona-Tailored Responses
We now compare Persona Tailoring (PT) to models trained on standard preference data [^8] Preference data is mostly for fine-tuning small, domain-specific models Ouyang et al. (2022), so we omit large, general models (e.g., ChatGPT) that we cannot feasibly fine-tune., showing PT improves personalization based on LLM (section 5.1, section 5.2, section 5.3) and eight users’ judgments (section 5.4, section 5.5).
@@ 5.1 Persona Tailoring Aids Personalization
We first confirm PT boosts personalization while maintaining quality versus models using standard preference datasets. PT always enhances personalization across generation strategies with varying resource demands—FS, SFT, and DPO—with minor response quality losses, shown via large $\Delta\textsc{PQ}$ (Table 2). Thus, practitioners seeking to improve personalization via prompting/training should capture why users prefer responses in data collection. Further, if preference data has already been curated, abduction via PI is a simple but effective augmentation strategy that largely improves personalization over diverse domains: dialogue, qa, and education.
@@ 5.2 $\textsc{PT}_{\textsc{dpo}}$ Supercharges Personalization
We verify $\textsc{PT}_{\textsc{dpo}}$ is the best method by comparing each PT model to FS without personas. Responses have large length differences, so to control for verbosity Zheng et al. (2024a), we compare model outputs with the same sentence count; the trend is similar without this check (Appendix A.7). Each method shows $\Delta\textsc{PQ}$ gains on Mnemonic, but SFT shows minor losses on BeaverTails (Table 3). Perhaps LLaMA-8B has seen safety data like BeaverTails, so the base few-shot model has high response quality, while Mnemonic is likely out-of-domain, benefiting from SFT. Regardless, $\textsc{PT}_{\textsc{dpo}}$ excels in $\Delta\textsc{PQ}$. Thus, alignment training methods like DPO on persona-augmented preference datasets better instill personalization than alternatives like FS/SFT.
@@
5.3 DPO Needs Personas for Personalization
Having seen $\textsc{PT}_{\textsc{dpo}}$’s benefits (section 5.2), we now test if personalization requires training on personas: can DPO trained without personas tailor to input personas in inference? Perhaps it is doable, as LLMs generalize to unseen instructions Deb et al. (2022), and personas are instructions. To answer this, we use BeaverTails/Anthropic which have first-person prompts, so we can add our personas to prompts by writing them in first person (“the user is X”$\rightarrow$“I am X”). We also test Mnemonic, but as prompts are vocabulary terms, models likely cannot generalize to personas; this is another benefit of our method, as we can improve personalization in any preference dataset. In the datasets, we compare outputs from $\textsc{PT}_{\textsc{dpo}}$ and DPO using input personas $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ via metrics from Prometheus (section 4.3). This ensures DPO and $\textsc{PT}_{\textsc{dpo}}$, which train on majority chosen outputs and thus may implicitly tailor to popular needs, can still aid less popular needs in $\mathcal{P}_{\texttt{R}}$ (section 3.3).
$\textsc{PT}_{\textsc{dpo}}$ nearly always bests DPO in personalization and quality, showing PT’s strength (Table 4). $\textsc{PT}_{\textsc{dpo}}$ also has more gains over DPO on $\mathcal{P}_{\texttt{R}}$ vs $\mathcal{P}_{\texttt{C}}$ (mean $\Delta\textsc{PQ}$ of 23.7 on $\mathcal{P}_{\texttt{R}}$ vs 13.4 on $\mathcal{P}_{\texttt{C}}$), so DPO sometimes adapts to needs in $\mathcal{P}_{\texttt{C}}$, but rarely uncommon ones in $\mathcal{P}_{\texttt{R}}$. Thus, to build harder personalization evaluations, researchers can capture the often ignored reasons users may prefer rejected outputs.
@@ 5.3 DPO Needs Personas for Personalization
Average DPO $\textsc{PT}_{\textsc{dpo}}$ 45.8/27.4/26.7 39.3/29.1/31.5 +18.4
@@
5.4 PT Personalizes to User-Specified Needs
To show PT aids real user needs, we recruit eight students of varied research backgrounds who use LLMs (Appendix A.12). We get twelve queries $q$ students may ask LLMs (e.g., job search) in BeaverTails and HHH. For both datasets, four users each write three personas $\mathcal{P}$ for each $q$ so models must adapt to $\mathcal{P}$, then rate $\textsc{PT}_{\textsc{dpo}}$ and DPO outputs (using $q$ and $\mathcal{P}$ as inputs) from 1–5 on Answerability (answering $q$) and Personalization (adapting to $\mathcal{P}$); this mirrors our judge evaluation (section 4.3) and ensures models support the needs of users who wrote them.
On BeaverTails, both models have high answerability but $\textsc{PT}_{\textsc{dpo}}$ is significantly more personalized (Figure 6). Anthropic HHH also shows minor improvements—with larger gains in personalization compared to degradation in answerability—reflecting our offline LLM-as-a-judge evaluation (section 5.1). Overall, $\textsc{PT}_{\textsc{dpo}}$ improves personalization, showing models trained on LLM-inferred personas can generalize to real user-specified needs. Since users perceive $\textsc{PT}_{\textsc{dpo}}$ as more helpful, future work can test how PT impacts long-term trust or engagement Serino et al. (2005) and if it aids downstream tasks Wang et al. (2024b); Mozannar et al. (2025).
Figure 6: Testing how well models aid user-specified needs. On BeaverTails, PT largely boosts personalization without losing answerability (Dror et al., 2018, 95% bootstrapped CIs).
5.5 Promises and Pitfalls of Persona Tailoring
We show strengths and issues of $\textsc{PT}_{\textsc{dpo}}$ in Figure 7. $\textsc{PT}_{\textsc{dpo}}$ alters answers for information needs and tailors mnemonics to learning styles (blue), showing its promise for downstream personalization tasks. However, $\textsc{PT}_{\textsc{dpo}}$ assumes all personas are harmless, leading to sycophancy Sharma et al. (2024); adversaries can exploit this and use personas to get inaccurate, biased, or irrelevant text (red). To solve this, we propose three safeguards for future work: curating undesired personas to teach LLMs when to abstain Wang et al. (2024c), system prompting to enable $\textsc{PT}_{\textsc{dpo}}$ to ignore harmful requests Zheng et al. (2024b), and flagging potentially harmful personas before $\textsc{PT}_{\textsc{dpo}}$ uses them Inan et al. (2023).